Line GH has endpoints G(-3,2) and H(3,-2). Find GH to the nearest tenth.
Found 2 solutions by stanbon, oberobic:
Answer by stanbon(57984)

Line GH has endpoints G(-3,2) and H(3,-2). Find GH to the nearest tenth.
-----
Assuming you are looking for the length of GH you get:
GH = sqrt[(3--3)^2 + (2--2)^2]
---
GH = sqrt[36+16]
---
GH = sqrt[52]
---
GH = 2sqrt(13) = 7.2 to the nearest tenth.
================
Cheers,
Stan H.

Answer by oberobic(2304)

The line GH from (-3,2) to (3,-2) can be depicted as the hypotenuse of a triangle.
The other corner of the triangle is at (-3,-2).
.
a = vertical from (-3,-2) to (-3,2). It has length 4.
.
b = horizontal from (-3,-2) to (3,-2). It has length 6.
.
c^2 = 4^2 + 6^2
.
c^2 = 16 + 36
.
c^2 = 54
.
c = sqrt(54)
.
sqrt(54) = sqrt(9*6)
.
sqrt(9*6) = 3*sqrt(6)
.
c is approximately = 7.3485
.
Done.